OpenStudy (anonymous):

Suppose you are an event coordinator for a large performance theater. One of the hottest new Broadway musicals has started to tour and your city is the first stop on the tour. You need to supply information about projected ticket sales to the box office manager. The box office manager uses this information to anticipate staffing needs until the tickets sell out. You provide the manager with a quadratic equation that models the expected number of ticket sales for each day x. ( is the day tickets go on sale). a. Does the graph of this equation open up or down? How did you determine t

OpenStudy (anonymous):

a. Does the graph of this equation open up or down? How did you determine this? b. Describe what happens to the tickets sales as time passes. c. Use the quadratic equation to determine the last day that tickets will be sold. Note. Write your answer in terms of the number of days after ticket sales begin. d. Will tickets peak or be at a low during the middle of the sale? How do you know? e. After how many days will the peak or low occur? f. How many tickets will be sold on the day when the peak or low occurs? g. What is the point of the vertex? How does this number relate to your answers in parts e. and f? h. How many solutions are there to the equation ? How do you know? i. What do the solutions represent? Is there a solution that does not make sense? If so, in what ways does the solution not make sense?

OpenStudy (anonymous):

i think its graph shoul be +ve

OpenStudy (anonymous):

think we are missing a piece of information yes?

OpenStudy (anonymous):

usually something like "for every dollar increase in price there are 3 less tickets sold" or something

OpenStudy (amistre64):

the numbers of tickets increase and then hit a plateu then begin to drop; opens down

OpenStudy (amistre64):

without the equation to go by; you cant really determine the rest of it

OpenStudy (anonymous):

because i am almost certain that you will have something like \[y=ax^2+bc+c\] with a negative number for a. then 1) the equation opens down 2) as time passes first then increase and then decrease 3) guess we need the equation for this one 4) ditto. need the vertex am i missing something??

OpenStudy (anonymous):

haven't given this award for a while

OpenStudy (anonymous):

Ticket= -0.2x^2+12x+11 is the equation that I didnt not get in there sorry

OpenStudy (anonymous):

aaaaaaaaaaaaaaaaaaaaaaaah

OpenStudy (anonymous):

now we have it all!

OpenStudy (anonymous):

\[T(x)=-0.2x^2+12x+11\]

OpenStudy (anonymous):

a) parabola faces down b) ticket sales first increase and then decrease as time goes on

OpenStudy (anonymous):

c) set = 0 and solve to get \[0= -0.2x^2+12x+11\] or \[-2x^2+120x+110=0\] or better yet \[x^2-60x-55=0\] and use the quadratic formula to solve for x.

OpenStudy (anonymous):

We look at the general quadratic \[q(x) = ax^{2} + bx + c,\]where a is nonzero. If a > 0, then the parabola is comcave upward, hence it has a global minimum at (-b/(2a), q(-b/(2a)).

OpenStudy (anonymous):

you should get something close to 61

OpenStudy (amistre64):

i came across this version of the quad formula \[\frac{-2c}{b\pm\sqrt{b^2-4ac}}\]

OpenStudy (anonymous):

If a < 0, the parabola is concave downward, then we have a global maximun at (-b/(2a), q(-b/(2a)).

OpenStudy (anonymous):

you will have peak sales at vertex of this parabola, which will occur at \[-\frac{b}{2a}\] where \[a=-.2,b=12\] so \[-\frac{b}{2a}=-\frac{12}{2\times -.2}=\frac{12}{.4}=\frac{120}{4}=30\]

OpenStudy (anonymous):

peak sales in 30 days. the number of sales will be whatever you get when you replace x by 30 in original expression

OpenStudy (anonymous):

the point of the vertex is \[(30,T(30))\] which represents the day of the peak sales (day 30) and the number of sales you get on that day (i will let you compute this)

OpenStudy (anonymous):

by the last question i assume it means how many solutions to the equation \[.2x^2+12x+11=0\] are there, and there are 2, but the negative one does not make sense in this context because you cannot go back in time. wish sometimes, but it is impossible

OpenStudy (anonymous):

and that is that

OpenStudy (anonymous):

so what is c,d,e,f,g,h then I am confused

OpenStudy (anonymous):

ss can u help more please

OpenStudy (anonymous):

How do u determine that the parabola faces down?

OpenStudy (anonymous):

If the highest term has a negative coefficient

OpenStudy (anonymous):

Will tickets peak or be at a low during the middle of the sale? How do you know?

OpenStudy (anonymous):

Look at the leading coefficient. _t's -0.2, a negative number, so the parabola is concave downward, so it has a peak, or global maximun.

OpenStudy (anonymous):

so that is the answer to d.?

OpenStudy (anonymous):

It will peak during the middle of the sale.

OpenStudy (anonymous):

how do we know?

OpenStudy (anonymous):

To get the peak value, a = -.2, b = 12, c = 11,- -b/(2a) = -(12/(-.4) = 30. The thirtieth day is when you have peak ticket sales. To determine how many tickets sold on that day, evaluate \[-0.2 x^{2}+ 12 x + 11\]at x = 30.

OpenStudy (anonymous):

On the thirtieth day of ticket sales, 3,431 tickets are sold.

OpenStudy (anonymous):

I am trying to finish d what was that?

OpenStudy (anonymous):

hero can you help me please?

OpenStudy (anonymous):

abtrehearn is right

OpenStudy (anonymous):

It will peak

OpenStudy (anonymous):

The answer to d is that ticket sales will peak during the middle of the sale. The reason is that the graph is concave downward (leading coefficient is negative), so it has a peak at the middle of the sale.

OpenStudy (anonymous):

ok I need to rest of them to please?

OpenStudy (anonymous):

abtrehearn also revealed that the peak will occur on the 13th day and the ticket sales will be 3,431.

OpenStudy (anonymous):

To work part e, get the vertex pf the parabola. Its x-coordinate is -b/(2a) = -12/(-.4) = -30. Peak ticket sales occur on the thirtieth day of the sale. That answers part e.

OpenStudy (anonymous):

Will tickets peak or be at a low during the middle of the sale? How do you know

OpenStudy (anonymous):

-12/(-.4 = 30, not -30.

OpenStudy (anonymous):

Peak sales occur on day 30 of the sale.

OpenStudy (anonymous):

how do we know?

OpenStudy (anonymous):

Do you renenber how to find the vertex of a parabola?

OpenStudy (anonymous):

How does this number relate to your answers in parts e. and f?

OpenStudy (anonymous):

It gives you when maximum ticket sales occur and how many tickets are sold that day.

OpenStudy (anonymous):

The x-coordinate is when maximum sales occur, and the y-coordinate of the vertex gives you how many tickets are sold that day.

OpenStudy (anonymous):

Takd a look at the graph of \[y = -0.2 x^{2} + 12x + 11.\]

OpenStudy (anonymous):

How many solutions are there to the equation ? How do you know?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Do you see where the peak is?

OpenStudy (anonymous):

There's most likely two solutions

OpenStudy (anonymous):

According to the graph that I seen there is no peak

OpenStudy (anonymous):

Did you make sure that there was a negative afront the highest coefficient?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I'm going to post what it should look like

OpenStudy (anonymous):

According to the function you posted....it has one peak and two solutions

OpenStudy (anonymous):

What do the solutions represent? Is there a solution that does not make sense? If so, in what ways does the solution not make sense?

OpenStudy (anonymous):

ok that helped some

OpenStudy (anonymous):

Any solution that is negative will not make sense

OpenStudy (anonymous):

what ways does it now make sense please

OpenStudy (anonymous):

there's no such thing as negative days

OpenStudy (anonymous):

First we need to find the solutions

OpenStudy (anonymous):

The solutions I have are (-0.903,0) and (60.9, 0)

OpenStudy (anonymous):

Basically the one that's negative obviously doesn't make sense

OpenStudy (anonymous):

The other one means that after 60 days, the sale is over.